OpenStudy (anonymous):

multiply 8 times the square root of 3 by 9 times the square root of 27.

OpenStudy (anonymous):

635.256

OpenStudy (slaaibak):

wait. Across I think you missed something. There by 8 * sqrt(3) * 27 * sqrt(3) You accidentally "factored" should have been 3 * 27 * 8 = 648

OpenStudy (slaaibak):

Or well, not factored. Just overlooked the sqrt(3)

OpenStudy (across):

You are right. I made a mistake: \[\sqrt{3}\cdot\sqrt{3}=3.\]Oh, my!